### v1

# Towards more realistic evaluation of LLM-based code generation: an experimental study and beyond

Abstract.
To evaluate the code generation capabilities of Large Language Models (LLMs) in complex real-world software development scenarios, many evaluation approaches have been developed. They typically leverage contextual code from the latest version of a project to facilitate LLMs in accurately generating the desired function. However, such evaluation approaches fail to consider the dynamic evolution of software projects over time, which we refer to as evolving-ignored situation, leading to issues of future context leakage and useful context missing. This in turn results in inaccurate evaluation of LLMs’ performance. In this paper, we conduct an empirical study to deeply understand LLMs’ code generation performance within settings that reflect the evolving nature of software development. To achieve this, we first construct an evolving-aware repository-level code generation dataset, namely HumanEvo, equipped with an automated execution-based evaluation tool. Second, we manually categorize HumanEvo according to dependency levels to more comprehensively analyze the model’s performance in generating functions with different dependency levels. Third, we conduct extensive experiments on HumanEvo with seven representative and diverse LLMs to verify the effectiveness of the proposed benchmark. We obtain many important findings through our experimental study. For example, we find that previous evolving-ignored evaluation approaches lead to inflated performance of the LLMs, ranging from 10.0% to 61.1%. Based on the findings, we give actionable suggestions on more realistic evaluation of LLMs on code generation. We also build a shared evolving-aware code generation toolbox to facilitate future research. Replication package including source code, datasets and appendix is available at https://github.com/DeepSoftwareAnalytics/EvoEval.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austria
1. Introduction
In recent years, the task of code generation based on LLMs has drawn widespread attention (Zheng et al., 2023b, a; Brown et al., 2020; Dong et al., 2023; Jiang et al., 2023; Zhang et al., 2023a; Hong et al., 2024; Qian et al., 2023). Many code LLMs (Rozière et al., 2024; Li et al., 2023; Guo et al., 2024; Zheng et al., 2023c; Chowdhery et al., 2023; Allal et al., 2023; Luo et al., 2024; Christopoulou et al., 2022; Shen et al., 2023; Nijkamp et al., 2023; Chen et al., 2021b; Li et al., 2022; Fried et al., 2023) are being employed to empower programming assistants, which are playing a vital role in practical software development.
Recently, several repository-level code generation benchmarks have emerged to better simulate real-world development scenarios, such as CoderEval (Yu et al., 2024), RepoBench (Liu et al., 2024), EvoCodeBench (Li et al., 2024a), and RepoEval (Zhang et al., 2023b). They sample functions from real-world projects as programming questions, aiming to reflect the performance of LLMs in actual development by prompting them to generate these functions according to the target function description and project context.
Figure 1. Issues of future context leakage and useful context missing
However, current benchmarks ignore the evolution nature (Voinea et al., 2005; Cleland, 2004; Robles et al., 2005) of real-world projects by treating the latest version of the repository as context, leading to inaccurate reflection of real development scenario. We identify two issues caused by this inaccurate reflection as depicted in Figure  1. The upper part of Figure 1 shows that, over time, the project repository has undergone multiple commits, including adding new functions and removing functions. In the lower part of Figure 1, we showcase the two issues that may arise from ignoring the evolution nature of a project, which we refer to as evolving-ignored situation. Example 1 illustrates the future context leakage issue that the evolving-ignored situation may lead to. Specifically, in the latest repository version, a code snippet committed in 2019 is highly similar to another snippet committed in 2022. Suppose the code snippet from 2019 is selected as the target function for LLMs. In that case, the code snippet from 2022 should not be included as project context during context acquisition even though they share high similarity. This is because the developer did not have access to the code committed in 2022 when writing the target function. Therefore, providing this code snippet as context to the model could pose the issue of future context leakage, resulting in an overestimated evaluation of LLMs’ code generation performance. Example 2 illustrates the useful context missing issue that the evolving-ignored situation may lead to. The target function is highly similar to the function in the pink box, and programmer deletes the function in the pink box after committing the target function. This suggests that when developers write the added code snippet, they can refer to that similar code snippet for inspiration. However, if the latest version of the project is utilized as context source, the model may fail to generate the target function correctly due to the absence of this useful context. This leads to an underestimation of LLMs’ code generation performance.
In this paper, we conduct an empirical study to comprehensively reveal LLMs’ repository-level code generation performance. To accomplish this goal, we construct HumanEvo, a novel evolving-aware repository-level code generation benchmark, which better simulates real-world development processes and reflects the evolving nature of projects over time. Specifically, projects evolve continuously throughout the development process, meaning that the project context available to programmers when writing different functions is constantly changing. To ensure that the context we provide precisely encompasses all the context available to the programmer when completing the target function (neither including additional code introduced by subsequent version changes nor omitting code that may have been updated or deleted due to version change), we roll back the entire project repository to the state before the target code was committed and evaluate the performance of LLMs on the rolled-back project.
We conduct a rigorous data collection pipeline to construct HumanEvo. We start with selecting high-quality projects and then collecting a large number of pull requests (PRs) from these projects. Then, we perform initial filtering on the crawled PRs to meet three attributes to guarantee the quality of the newly added functions in PRs. Additionally, to ensure that the newly added functions are covered by the test suite, we establish a runtime environment for each PR and run the corresponding project’s testing framework. After all this, we obtain 400 programming questions, 200 for Python and 200 for Java. In our benchmark construction process, our selection of functions revolves around PRs rather than commits. First, functions in a merged PR are typically reviewed by project managers (Rahman and Roy, 2014a), ensuring the quality of the added functions. Second, a PR usually incorporates test code for the added functions, guaranteeing their coverage by the project’s testing framework.
We conduct an extensive evaluation of multiple mainstream LLMs (the CodeLlama series (Rozière et al., 2024), DeepSeekCoder series (Guo et al., 2024), GPT-4 (OpenAI, 2024),GPT-3.5-Turbo) on HumanEvo. Experimental results reveal the performance of LLMs in more practical development scenarios and emphasize the importance of evolving-awareness, showcasing how LLMs gradually deviate from reality in evolving-ignored situations. Specifically, we first compare the performance of LLMs under evolving-ignored and evolving-aware settings, and experimental results indicate that LLMs exhibit inflated performance under the evolving-ignored setting. To investigate whether both issues depicted in Figure 1 have certain impacts on the models, we further analyze task instances successfully completed by the models under both settings and find that both “future context leakage” and “useful context missing” have significant impacts on the models. According to the empirical results, we highlight four key findings as follows:
  1. (1)
    Previous evolving-ignored evaluation approaches lead to inflated performance of the LLMs, ranging from 10.0% to 61.1%, revealing how previous evaluation methods provided the models with a deviated programming scenario from reality.
  2. (2)
    LLMs’ success rates in generating functions with intra-class dependencies decreased by approximately 14.0% on average, and those with inter-class dependencies by approximately 30.9%, under the evolving-aware setting compared to the evolving-ignored setting. This suggests that functions with broader dependencies are more impacted on different evaluation settings.
  3. (3)
    Under fine-granularity evolving-ignored settings, the discrepancy between model performance and real-world effectiveness increases over time, exhibiting a progressive, upward trend in inflated evaluation metrics as projects evolve.
  4. (4)
    Detailed docstrings significantly enhance the LLMs’ ability to accurately generate the expected code.
In summary, our contributions can be summarized as follows:
  • We have identified a common flaw in previous repository-level code generation benchmarks: they are all evolving-ignored, causing inaccurate simulations of real-world development scenarios.
  • We introduce a new benchmark HumanEvo, an evolving-aware repository-level code generation benchmark that addresses the evolving-ignored issue, to conduct a comprehensive empirical study on the performance of LLMs in repository-level code generation.
  • Through extensive experimentation, we substantiate the evolving-ignored situation leads to inflated performance of the models, with LLMs showing a performance inflation ranging from 10.0% to 61.1%.
  • For convenient usage of HumanEvo, we release all data, code, and the evaluation platform based on Docker image that provides runtime environments for all projects, facilitating the automated execution of the test suites.
2. Background
2.1. Code Generation Benchmark
Previously, the vast majority of benchmarks were constructed to evaluate the performance of LLMs in generating standalone functions(Athiwaratkun et al., 2023; Cassano et al., 2022a; Jain et al., 2022; Zan et al., 2022b; Du et al., 2023; Ding et al., 2024; Khan et al., 2023), with the most representative ones being HumanEval  (Chen et al., 2021a) and MBPP  (Austin et al., 2021). HumanEval comprises 164 manually crafted programming problems covering language comprehension, reasoning, algorithms, and simple mathematics. MBPP consists of 974 short Python function problems primarily designed for novice programmers.
Finding previous benchmarks fail to reflect complex programming scenarios, researchers propose CoderEval (Yu et al., 2024), the first repository-level code generation benchmark, to evaluate the performance of existing LLMs in real-world development by prompting them to generate functions selected in actual complex software. CoderEval meticulously selects 230 Python and 230 Java programming tasks from popular real-world open-source projects, evaluating the functional correctness of generated code based on execution. To better simulate real-world development scenarios, EvoCodeBench (Li et al., 2024a) first analyses the distribution of standalone and non-standalone functions in a large number of real-world projects, then constructs a benchmark with a dependency distribution more in line with real-world scenarios.
Although the aforementioned works prompt LLMs to generate functions selected from real software projects to evaluate LLM performance in real development scenarios, we recognize a common flaw in all these benchmarks: they ignore the evolution nature of project repositories. This implies that when we prompt LLMs to generate different functions, the project context provided to LLMs should also consider the evolution of the project. Taking the example of software development processes based on GitHub project management, when writing functions, the project context that programmers can refer to is the entire project repository up to the commit preceding the target functions. Therefore, when prompting LLM to generate target functions, we should roll back the entire project repository to the state before the target function was committed.
2.2. Repository-level Code Generation
Many technical studies have been proposed to better tackle repository-level code generation tasks (Shrivastava et al., 2024; Zhou et al., 2023; Zhang et al., 2023b; Zan et al., 2023, 2022a; Bairi et al., 2023; Liang et al., 2024; Eghbali and Pradel, 2024; Zhang et al., 2024; Phan et al., 2024; Wang et al., 2024). DocPrompting (Zhou et al., 2023) uses natural language (NL) intent to retrieve documents related to the target code and then sends NL intent and retrieved documents to LLMs. CoCoMIC (Ding et al., [n. d.]) develops a cross-file context finder, which is a static code analysis tool used to retrieve the most related cross-file context. MGD (Agrawal et al., 2024) leverages IDEs to assist in providing context. IDE is invoked to provide useful suggestions, when reaching the pre-defined trigger point. RLPG (Shrivastava et al., 2023) analyses dependencies potentially existing in the target function, then extracts project context that may be conducive to generating the target function. RepoCoder (Zhang et al., 2023b) first slices the source code files in the repository to construct a retrieval library, then uses the generated problem as a query, calculates similarity to obtain project context, and feeds the project context and problem to the model to generate initial results. After obtaining the initial results, the previous generation results are used as queries for further querying to obtain more context. This process is repeated to continuously optimize the generation results. $A^{3}$-CodGen (Liao et al., 2024) first constructs a repository knowledge base covering knowledge of third-party libraries and extracts information on mobile and code context. During the code generation process, LLM first generates an initial version of code based on NL intent, then uses NL intent and the initial version of code to retrieve as much foundational knowledge as possible from the knowledge base for the LLM. Finally, local, global, and third-party library information is integrated into the prompt to allow the model to generate the final result.
However, all the mentioned works treat the latest version of the project repository as the context source, falling into the category of evolving-ignored scenarios. Although in RepoCoder, researchers begin to take notice of this issue by considering all code after the target function in the corresponding file as “future code”, this still falls short of reality and introduces a lack of the context that should be present. Therefore, we propose a novel repository-level code generation benchmark, HumanEvo, an evolving-aware pragmatic benchmark, to better simulate real development scenarios for LLMs.
2.3. LLMs for Code Generation
CodeLlama (Rozière et al., 2024) is further trained based on Llama2 (Touvron et al., 2023). It is capable of understanding and generating code in various programming languages, including but not limited to Python, Java, JavaScript, etc. CodeLlama not only generates complete code but also provides suggestions and optimization solutions for specific code snippets, greatly assisting developers. During the training of the base model, weights are initialized with an equivalent parameter amount of the Llama2 model, followed by training on a dataset of 500 billion tokens of code. StarCoder (Li et al., 2023) is derived from StarCoderBase, which is trained using 10 trillion tokens from The Stack, a vast collection of licensed GitHub repositories with inspection tools and selection exit procedures. The DeepSeek-Coder (Guo et al., 2024) takes into account the often intricate dependency relationships inherent in real-world projects. The model needs to fully consider these dependencies before completing programming tasks. Therefore, to equip DeepSeek-Coder to handle real-world complex projects, the training data is structured to support repository-level comprehension, enhancing the model’s understanding of complex projects. This enables the model to utilize cross-file context effectively. MetaGPT (Hong et al., 2024) constructs a multi-agent collaborative framework, employing Standardized Operating Procedures and involving multiple roles such as programmers and product managers to collaborate on project completion. The results generated by agents are not saved in conversational form but in structured output formats such as charts, files, etc. Communication between agents is also conducted through structured files. MetaGPT also implements an iterative development pipeline, optimizing results through multiple rounds of iteration by recording information generated during the development process by different roles. Similarly, ChatDev (Qian et al., 2023) adopts a multi-agent framework, allocating and assigning different AI agents to various job functions. ChatDev facilitates communication among different role agents, forming a complete software development process where each agent performs its respective tasks while collaborating uniformly. In addition to these models, there are many other excellent models used in the field of code generation (Luo et al., 2024; Christopoulou et al., 2022; Shen et al., 2023; Nijkamp et al., 2023; Chen et al., 2021b; Li et al., 2022; Fried et al., 2023).
2.4. Project Evolution
GitHub-hosted projects undergo continuous evolution over time (Negoita et al., 2019). GitHub provides an open platform where the development progress, issues, and solutions of projects are visible to everyone (Dabbish et al., 2012; Kalliamvakou et al., 2015). When users identify missing features in a project, they typically submit an issue to describe their needs to the developers (Kallis et al., 2021). To address the issue, developers will create a new branch and write code to implement the new feature or fix the bug according to user requirements. Once the code changes are complete, developers commit them to the new branch, push the branch to the GitHub repository, and then initiate a pull request to propose merging these code changes into the main branch (Rahman and Roy, 2014b). Finally, project administrators will review the changes in the pull request and, upon approval, merge them into the main branch.
3. HumanEvo Benchmark
In this section, we provide a detailed introduction to HumanEvo from three perspectives: benchmark overview, benchmark construction pipeline, and benchmark characteristics.
3.1. Benchmark Overview
Figure 2. Task instance overview.
HumanEvo consists of a total of 400 task instances, comprising 200 Python programming questions and 200 Java programming questions, respectively. As illustrated in Figure 2, each task instance is associated with a unique identification number. The Owner/Repo field helps identify the GitHub project from which the selected function originates, while the Version field indicates the version of the project during the implementation of this function. Base commit refers to the last commit on the branch before the target function is committed. We use it to restore the project repository to the state before the target function was implemented. The Function declaration of the target function will be provided to the LLM as part of the prompt, along with its corresponding docstring. Since the docstring styles of functions in pragmatic projects vary, with some adopting line-by-line comments and others even lacking corresponding comments, we think it is necessary to rewrite the docstring for each selected function. As depicted in Figure 2, we provide two styles of docstrings. The Brief docstring provides a concise summary of the functionality of the target function, while the Detailed docstring provides a thorough description of the function’s functionality, inputs, and outputs.
3.2. Benchmark Collection Pipeline
Figure 3. HumanEvo construction pipeline.
As depicted in Figure 3, the construction pipeline of HumanEvo consists of six steps, mainly revolving around the collection of high-quality PRs. Here are the reasons why we construct HumanEvo through PRs. We use PRs to construct HumanEvo primarily for the following three reasons: First, since prompting LLMs to generate repository-level functions can be viewed as an incremental development process, and PRs for adding new features exactly mirror the incremental development process in the real world, we choose to have LLMs generate the newly added functions in PRs to simulate real-world development scenarios. Second, the quality of functions in PRs is guaranteed. Before merging the newly added code in PRs into the project, repository administrators conduct a review on it, which allows us to select high-quality task instances directly. Third, the base commit in PRs helps us roll back the project repository to the state before the target function was committed. Finally, by processing PRs appropriately, we can further determine whether the target functions are covered by the project’s testing framework.
3.2.1. Projects selection
In order to make HumanEvo more representative and adequate, we adopt a comprehensive approach to selecting functions from various open-source projects. For Python, we select high-quality real-world projects from a curated list of the top 5,000 most downloaded PyPI libraries during 2023111https://hugovk.github.io/top-pypi-packages/. Subsequently, we deploy the selected projects and run their testing frameworks to ensure each of them can run correctly and stably. We then proceed to collect PRs for these repositories through the GitHub developer API [^2] https://api.github.com. For Java, as widely recognized projects typically indicate extensive documentation, structured open-source development guidelines, and functional, well-formatted code, our approach involves searching for the top 200 starred projects on GitHub. Then, we deploy every project on the list to select projects that can be successfully deployed and pass the corresponding test suite. This methodology ensures that the functions selected for evaluation in HumanEvo are representative of real-world programming scenarios, encompassing a diverse range of projects and languages.
3.2.2. PR crawling
We crawl a large number of initial PRs from the popular open-source repositories selected in the procedure above. Specifically, we obtain 12,039 PRs and meticulously record metadata pertaining to these PRs for subsequent filtering processes.
3.2.3. PR attribute-based filtering
In this step, we preliminarily filter the crawled PRs based on the following three attributes.
  • The status of the PR is “Merged”. A “Merged” status indicates that the code changes associated with the PR were accepted and incorporated into its parent repository.
  • The PR must introduce new functions. These newly added functions are typically intended to develop the software project by adding new features, improvements, or fulfilling general development objectives, which can be selected as programming questions.
  • The PR must introduce at least one or more new tests, and these tests must cover the selected function.
All these attributes aim to ensure that the subsequently selected functions are of high quality, have undergone rigorous review and are covered by the project’s testing framework.
3.2.4. PR execution-based filtering
In this step, we perform execution-based filtering for the PRs filtered in the attributes filtering step.
Firstly, we prepare the execution environment for each candidate task instance in HumanEvo. For each task instance, we first determine the version of the project corresponding to the PR, then manually identify the required dependency environment for running the project at that version. For example, for a Python task instance, we confirm the required Python version and the types and versions of third-party libraries it depends on. After obtaining this dependency environment information, we create a virtual execution environment for each version of the project. Task instances that can not install the dependency environment properly are filtered out during this step. After creating the execution environment, we execute the project’s testing framework in the corresponding environment.
Secondly, for each selected function, we execute the project’s test framework in its required environment to ensure that the selected function is covered by the project’s test framework. Specifically, we split the PR’s patch file, which is also known as the code diff file, into two parts, namely the production code patch and the test code patch. For each candidate task, we apply the PR’s test patch and log the associated test results before and after applying the PR’s production patch. Then, we filter out task instances with no test where its status changes from a fail to pass, which indicates the newly added functions are covered by the test suite. After filtering out instances without a fail-to-pass transition, we obtain 400 PRs.
3.2.5. Function selection
In this step, we manually select functions from the PRs that have passed the filtering process above to serve as programming questions. After completing the above filtering, we confirm that the remaining PRs meet our requirements. Therefore, we select newly added functions from these PRs to as programming questions. During the selection process, to avoid choosing basic functions, we filter out initialization functions, constructors, and destructors, to ensure the quality of the selected programming questions. Finally, we acquire a total of 400 programming questions.
3.2.6. Human annotation
In this step, we recruit six annotators with more than five years of programming experience to manually rewrite two different styles of docstrings for the selected functions and label each function with a dependency level.
We rewrite the docstring for each instance in HumanEvo to ensure the quality of the benchmark dataset and mitigate data leakage issues as much as possible. Specifically, in the process of data collection, we find that the docstring styles of actual projects can be divided into two categories: the detail and the brief, so in order to make our benchmark more diverse and better reflect the actual development scenarios, we provide two types of the docstring for each task instance. The detailed docstring describes the function of the target function in detail, as well as the types and meanings of its inputs and outputs. In contrast, a brief docstring only briefly describes what the target function does. After the annotators complete writing the docstrings, we conduct a cross-validation review, requiring each task instance to be verified and confirmed by two developers to ensure accuracy. If the annotators disagree with a docstring written by the previous annotator during the review process, they will discuss the issue together until all annotators reach an agreement.
In order to facilitate the evaluation of the ability of LLMs in generating functions with different dependency levels, we further categorize all task instances in HumanEvo into three categories: standalone, intra-class, and inter-class functions. A dependency denotes the invocation of elements defined within projects, and the dependency level is defined based on the position of the target function’s dependencies. Standalone functions refer to those that invoke or access only built-in functions and standard libraries. Intra-class functions rely on other functions within the same class. Inter-class functions involve calling functions within the same file or functions in other files.
3.3. Benchmark Characteristics
3.3.1. Evolving-aware benchmark
HumanEvo is the pioneering repository-level code generation benchmark to introduce the evolving-aware concept. We highlight a significant oversight in prior research: the developmental trajectory of a project is an evolving process, altering the project context available to programmers as the project advances through various stages. Previous studies have overlooked this temporal dimension, resulting in LLMs being fed with code that had not yet been implemented in the project when certain functions were developed. Consequently, it also misses some context that should have been present but might have been deleted or updated during the project’s evolution. This undoubtedly leads to the model facing an unrealistic development scenario. As a result, the performance of LLMs in real-world software development tasks remains unknown. Therefore, we are committed to enhancing HumanEvo to serve as a benchmark that better simulates real-world software code development scenarios, aiming to more accurately reflect the true performance of LLMs applied to pragmatic development tasks.
3.3.2. Strict data filtering process
We ensure the quality of the selected function from several aspects. First, we ensure that the selected projects are of high quality because high-quality projects generally have excellent maintenance, comprehensive testing frameworks, and good coding styles. For Python, we choose the projects with the highest download count in PyPI. For Java, we select projects from the top 200 projects on GitHub based on the number of stars, with an average of around 24k stars. Previous benchmarks have ensured the functionality of target functions by writing unit tests for each function separately. However, we believe that in a project with complex dependencies, adding a new function may have certain impacts on other parts of the project. Relying solely on unit tests for the target function is insufficient to ensure that the project can still function properly and pass the project’s test framework after inserting the function. Therefore, in this work, we directly run the project’s test framework to ensure the correctness of the target function’s functionality while ensuring that inserting the target function does not have a negative impact on other parts of the project. Of course, to ensure that the test framework covers the target function, during the function selection process, we verify that our selected functions are covered by the project’s test framework through execution-based validation. In other words, a “fail-to-pass” test result implies that the test is set up for the target function and the target function has passed the test.
Table 1. Dependency level in HumanEvo.
Dependency Level HumanEvo-Python HumanEvo-Java
Standalone 46 (23.0%) 26 (13.0%)
Intra-class 61 (30.5%) 61 (30.5%)
Inter-class 103 (51.5%) 113 (56.5%)
3.3.3. Human annotation
We manually rewrite two styles of docstrings for all task instances and label each of them with a dependency level. During the dataset construction process, we observe different docstring styles among these high-quality projects, which can be generally categorized into the detailed and the brief. Detailed docstring typically provides comprehensive descriptions of a function’s functionality, elucidates input and output parameter types, and annotates potential error scenarios. Conversely, a brief docstring succinctly outlines the function’s functionality in a few sentences. To enhance the diversity of our benchmark and better align with real-world development scenarios, we manually crafted two types of docstring for each function: detailed and brief. This provides users with more options for subsequent usage.
Furthermore, to facilitate the assessment of LLMs’ capability in generating functions with varying levels of dependency, we categorize the selected functions based on their dependency levels, including standalone, intra-class, and inter-class. Standalone functions can be implemented without relying on other functions within the project. Intra-class functions require dependencies on other functions within the same class. Inter-class functions imply more complex dependencies, with most of the dependent functions being located in other files within the project. Manual categorization of dependency levels aids in analyzing the performance of LLMs in generating functions of varying complexity, facilitating researchers in improving the application of LLMs in practical development scenarios based on different contexts. Moreover, as indicated in Table 1, the majority of our benchmark consists of functions with deep dependency levels. This suggests that our benchmark presents a sufficiently challenging scenario.
4. Experimental Design
In this section, we introduce the models used in the experiments, outline the research questions that the experiments aim to address and provide a detailed description of the experimental setups.
4.1. Model Selection
We categorize mainstream LLMs for code generation into two groups: open-source LLMs and closed-source LLMs. For open-source LLMs, we select the CodeLlama series (including CodeLlama-7B, CodeLlama-13B, and CodeLlama-34B) and the DeepSeekCoder series (including DeepSeekCoder-6.7B and DeepSeekCoder-33B). The CodeLlama and DeepSeekCoder series are popular models that perform well in code generation. For closed-source LLMs, we choose the commonly used commercial models: GPT-3.5-Turbo and GPT-4, with GPT-4 having demonstrated superior performance levels in previous benchmarks.
4.2. Research Questions
Our experiments intend to answer the following research questions:
  • RQ1: How effective is HumanEvo in benchmarking the performance of LLMs in repository-level code generation task?
    We conduct experiments under two different settings, evolving-aware and evolving-ignored, to explore the differences in model performance between these two settings and reveal the actual performance of the models in real development scenarios.
  • RQ2: How does the code generation performance change as repository evolves?
    We selected multiple versions of the project after the target function was committed as the context source to explore how the model’s performance gradually deviates from reality over time under the evolving-ignored setting.
  • RQ3: How different code descriptions in the dataset affect code generation performance?
    We conduct experiments to compare the performance of the model under different docstring styles.
4.3. Experimental Settings
Due to the large size of modern software repositories, it is not feasible to feed the entire repository to the model. Thus, previous researchers mainly obtain project context that might be beneficial for generating target functions through two approaches: retrieval-based and static analysis-based methods. In the experiments below, we cover both of these technical routes to demonstrate that the evolving-ignored situation leads to the deviation of the context source from reality.
4.3.1. Retrieval based project context acquisition
We adopt the standard retrieval-augmented-generation pipeline for repository-level code generation (Zhang et al., 2023b; Formal et al., 2022; Santhanam et al., 2022; Xiong et al., 2021). In the evolving-aware setting, we restore the project repository to the state before the target code was committed for each task instance. On the contrary, in the evolving-ignored settings, we directly utilize the latest project version as the context source. Then, we construct a corresponding context retrieval database by partitioning the source code files from the repository into a collection of code snippets. Subsequently, we utilize the docstring of the target code as the query for retrieval and compute the Jaccard similarity between the query and all code snippets in the context retrieval database. After sorting all the code snippets by similarity score, we select project contexts with higher similarity scores as project context for target function generation.
4.3.2. Static analysis based project context acquisition
Similar to the static analysis based repository-level code generation work (Shrivastava et al., 2023), we obtain code files in the repository that might help generate the target code through static analysis. First, we roll back the project repository to the corresponding state according to evolving-ignored or evolving-aware settings. Then, we extract the source code from these files as project context to feed to the model. There are four types of context sources: local file (the file where the target code resides), import file (files imported by the local file in the project repository), sibling file (files located in the same folder as the local file), and similar file (files in the project repository with names similar to the local file). We utilize four prompting strategies from (Shrivastava et al., 2023): local – extracting context solely from the local file; local+import – extracting code from both local and import files; local+sibling - extracting code from both local and sibling files; local+sim - extracting code from both local and similar files.
4.3.3. Evaluation metric
We assess the functionality correctness of code generated by LLMs with execution-based evaluation and utilize the commonly used metric Pass@1 for evaluation. Note that, to mitigate issues stemming from model-generated randomness, the experimental results presented subsequently are obtained by conducting three repeated experiments and averaging the results.
Table 2. Retrieval-based project context acquisition approach: model performance under evolving-aware (HumanEvo) and evolving-ignored (EI) settings. DepSCode is short for DeepSeekCoder.
Model Python Java
EI HumanEvo EI HumanEvo
CodeLlama-7B 29.0% 22.0% ($\downarrow 24.1\%$) 8.5% 6.0% ($\downarrow 29.4\%$)
CodeLlama-13B 30.5% 23.0% ($\downarrow 24.6\%$) 9.5% 7.0% ($\downarrow 26.3\%$)
CodeLlama-34B 31.5% 24.0% ($\downarrow 23.8\%$) 14.5% 10.0% ($\downarrow 31.0\%$)
DepSCode-6.7B 30.0% 23.0% ($\downarrow 23.3\%$) 13.0% 9.0% ($\downarrow 30.8\%$)
DepSCode-33B 31.0% 25.0% ($\downarrow 19.4\%$) 15.5% 11.5%( $\downarrow 25.8\%$)
GPT-3.5-Turbo 32.5% 25.5% ($\downarrow 21.5\%$) 18.0% 13.0% ($\downarrow 27.8\%$)
GPT-4 34.5% 26.5% ($\downarrow 23.2\%$) 20.5% 14.5% ($\downarrow 29.3\%$)
5. Evaluation Results
Table 3. Static analysis based project context acquisition approaches: model performance under evolving-aware (HumanEvo) and evolving-ignored (EI) settings.
Data Model local local+import local+sibling local+sim
EI HumanEvo EI HumanEvo EI HumanEvo EI HumanEvo
HumanEvo-Python CodeLlama-7B 27.0% 24.0%($\downarrow 11.1\%$) 28.0% 25.0%($\downarrow 10.7\%$) 28.5% 25.0%($\downarrow 12.3\%$) 29.0% 25.0%($\downarrow 13.8\%$)
CodeLlama-13B 28.5% 24.5%($\downarrow 14.0\%$) 29.5% 26.5%($\downarrow 10.2\%$) 29.5% 25.5%($\downarrow 13.6\%$) 30.5% 25.5%($\downarrow 16.4\%$)
CodeLlama-34B 29.5% 25.0%($\downarrow 15.3\%$) 30.0% 27.0%($\downarrow 10.0\%$) 30.0% 26.5%($\downarrow 11.7\%$) 30.5% 26.5%($\downarrow 13.1\%$)
DeepSeekCoder-6.7B 29.0% 24.0%($\downarrow 17.2\%$) 29.0% 24.5%($\downarrow 15.5\%$) 29.0% 25.5%($\downarrow 12.1\%$) 29.5% 26.0%($\downarrow 12.1\%$)
DeepSeekCoder-33B 30.5% 25.5%($\downarrow 16.4\%$) 31.0% 26.5%($\downarrow 14.5\%$) 30.5% 26.5%($\downarrow 13.1\%$) 31.0% 26.5%($\downarrow 14.5\%$)
GPT-3.5-Turbo 33.0% 26.5%($\downarrow 19.7\%$) 34.0% 27.0%($\downarrow 20.6\%$) 32.5% 26.5%($\downarrow 18.5\%$) 33.5% 27.5%($\downarrow 17.9\%$)
GPT-4 34.5% 27.0%($\downarrow 21.7\%$) 35.0% 28.0%($\downarrow 20.0\%$) 34.5% 27.5%($\downarrow 20.3\%$) 35.5% 28.0%($\downarrow 21.1\%$)
HumanEvo-Java CodeLlama-7B 16.5% 8.0%($\downarrow 51.5\%$) 15.0% 6.5%($\downarrow 56.7\%$) 12.0% 6.0%($\downarrow 50.0\%$) 13.5% 6.5%($\downarrow 51.9\%$)
CodeLlama-13B 15.0% 6.5%($\downarrow 56.7\%$) 16.0% 7.0%($\downarrow 56.3\%$) 15.5% 6.5%($\downarrow 58.1\%$) 15.5% 6.5%($\downarrow 58.1\%$)
CodeLlama-34B 17.5% 8.0%($\downarrow 54.3\%$) 14.5% 7.0%($\downarrow 51.7\%$) 12.5% 6.5%($\downarrow 48.0\%$) 14.5% 8.0%($\downarrow 44.8\%$)
DeepSeekCoder-6.7B 17.0% 8.5%($\downarrow 50.0\%$) 16.5% 10.0%($\downarrow 39.3\%$) 14.5% 9.0%($\downarrow 37.9\%$) 18.0% 7.0%($\downarrow 61.1\%$)
DeepSeekCoder-33B 16.5% 10.0%($\downarrow 39.3\%$) 16.5% 10.0%($\downarrow 39.3\%$) 18.5% 8.5%($\downarrow 54.1\%$) 19.0% 10.0%($\downarrow 47.4\%$)
GPT-3.5-Turbo 19.5% 11.5%($\downarrow 41.0\%$) 21.0% 14.0%($\downarrow 33.3\%$) 20.0% 12.0%($\downarrow 40.0\%$) 21.5% 13.0%($\downarrow 39.5\%$)
GPT-4 22.0% 13.5%($\downarrow 38.6\%$) 23.0% 14.5%($\downarrow 37.0\%$) 22.5% 14.0%($\downarrow 37.8\%$) 23.5% 14.5%($\downarrow 38.3\%$)
5.1. RQ1: How effective is HumanEvo in benchmarking the performance of LLMs in repository-level code generation task?
5.1.1. Overall performance
Table 2 and Table 3 present the performance of seven mainstream LLMs under two different experimental settings: Evolving-Ignored (where the context source is the latest version of the project repository) and Evolving-Aware (HumanEvo, where the context source is the project repository before the target function was committed). Table 2 shows the performance of retrieval based project context acquisition approach, and Table 3 shows the performance of static analysis based project context acquisition approaches.
All the experimental results shown in Table 2 and Table 3 show a significant decrease in performance under the evolving-aware setting compared to the evolving-ignored setting. This suggests that in the evolving-ignored setting, the context source provided to the model contains too much future context, leading to an unrealistically inflated performance of the LLMs.
5.1.2. Break down analysis
Table 4. Break down analysis. EA and EI stand for evolving-aware and evolving-ignored, respectively. ✓means success,
$\times$
meas failure.
Model Python Java
EA✓ EA✓ EA
$\times$
EA✓ EA✓ EA
$\times$
EI
$\times$
EI✓ EI✓ EI
$\times$
EI✓ EI✓
CodeLlama-7B 7 37 21 3 9 8
CodeLlama-13B 4 42 19 3 11 8
CodeLlama-34B 2 46 17 6 14 15
DeepSeekCoder-6.7B 4 42 18 7 11 15
DeepSeekCoder-33B 4 46 16 5 18 13
GPT-3.5-Turbo 2 49 16 10 16 20
GPT-4 0 53 15 4 25 16
We further analyze the generation results of LLMs to investigate whether both issues described in Figure 1 have certain impact on the final results. Table 4 illustrates the overlap of successful task instances between evolving-aware setting and evolving-ignored setting. Since the results from retrieval-based project context acquisition approach and static analysis based project context acquisition approaches are similar, subsequent analysis primarily focuses on the experimental results of the retrieval-based approach. As introduced above, the future context leakage issue will lead to inflated performance, while the useful context leakage issues issue will lead to underestimated performance. However, the results in Table 2 reflect that models’ overall performance only exhibits an inflated trend, driving us to delve deeper into whether the situation of useful context leakage has occurred. We investigate the overlap between task instances that LLMs successfully completed under evolving-ignored and evolving-aware settings.
In Table 4, EA stands for the evolving-aware setting and EI stands for the evolving-ignored setting. A ✓means successful, while a $\times$ means failure. So, EA✓EI $\times$ refers to the number of instances that pass under the evolving-aware setting but fail under the evolving-ignored setting, which is typically caused by useful context missing. EA $\times$ EI✓represents instances that fail under EA but pass under EI, which is typically caused by future context leakage. EA✓EI✓denotes the number of instances that pass under both settings. From the experimental results, it is evident that except for GPT-4 in HumanEvo-Python, which has a value of 0 for EA✓EI $\times$ situation, all other LLMs exhibit deviations from reality caused by the absence of necessary context in a certain quantity, showing that the issue of useful context missing have impact on evolving-ignored setting. However, due to the significantly higher number in EA $\times$ EI✓, the primary influencing factor remains the future context leakage, leading to exaggerated performance of LLMs.
Figure 4. LLM’s performance when using different project version as context source.
Table 5. Model performance in generating functions with deeper dependency levels.
Model Python Java
EI HumanEvo EI HumanEvo
Intra-Class CodeLlama-7B 32.8% 31.1% ($\downarrow 5.0\%$) 1.6% 1.6%
CodeLlama-13B 29.5% 27.9% ($\downarrow 5.6\%$) 2.9% 1.9% ($\downarrow 33.3\%$)
CodeLlama-34B 34.4% 27.9% ($\downarrow 19.0\%$) 8.2% 6.6% ($\downarrow 20.0\%$)
DeepSeekCoder-6.7B 34.4% 27.9% ($\downarrow 19.0\%$) 4.9% 3.3% ($\downarrow 33.3\%$)
DeepSeekCoder-33B 37.7% 34.4% ($\downarrow 8.7\%$) 3.3% 3.3%
GPT-3.5-Turbo 36.1% 29.5% ($\downarrow 18.2\%$) 4.9% 4.9%
GPT-4 37.7% 32.9% ($\downarrow 13.0\%$) 8.2% 4.9% ($\downarrow 40.0\%$)
Inter-Class CodeLlama-7B 27.2% 17.5% ($\downarrow 35.7\%$) 10.6% 5.3% ($\downarrow 50.0\%$)
CodeLlama-13B 31.1% 20.3% ($\downarrow 34.4\%$) 10.6% 6.2% ($\downarrow 41.7\%$)
CodeLlama-34B 30.1% 22.3% ($\downarrow 25.8\%$) 15.0% 10.6% ($\downarrow 29.4\%$)
DeepSeekCoder-6.7B 27.2% 20.4% ($\downarrow 25.0\%$) 15.9% 11.5% ($\downarrow 27.8\%$)
DeepSeekCoder-33B 28.2% 21.4% ($\downarrow 24.1\%$) 17.7% 13.3% ($\downarrow 25\%$)
GPT-3.5-Turbo 31.1% 23.3% ($\downarrow 25.0\%$) 21.2% 14.2% ($\downarrow 33.3\%$)
GPT-4 32.0% 23.3% ($\downarrow 27.3\%$) 22.1% 15.9% ($\downarrow 28.0\%$)
5.1.3. Dependency level
Functions with different dependency levels vary in their reliance on project context, and the evolving-ignored setting directly affects project context. Therefore, in this experiment, We examine the performance differences of LLMs in generating functions with deeper dependency levels, focusing on non-standalone functions in HumanEvo, which encompass intra-class and inter-class functions, across two settings: evolving-ignored and evolving-aware.
As shown in Table 5, under the evolving-aware setting, the accuracy of LLM in generating both intra-class and inter-class functions has noticeably declined. The decline is more pronounced for inter-class functions, as these functions typically require cross-file context, which is often more scattered. As the project evolves, its various parts tend to become more comprehensive. Therefore, under the evolving-ignored setting, the models can reference an increasing amount of context. Consequently, when the evolving-aware setting loses access to these contexts, the model’s performance significantly deteriorates.
RQ1 Summary: Experiments show an inflated trend in LLM’s performance under the Evolving-ignored setting. Upon further investigation, we discovered that both the presence of too much extraneous context and the disappearance of context that should exist have significantly affected the model’s performance, resulting in the model’s performance deviating from reality.
5.2. RQ2: How does the code generation performance change as repository evolves?
To further illustrate the impact of the evolving-ignored setting on the performance of LLMs as projects evolve, we conduct more in-depth experimental exploration. Typically, a project’s version number consists of three digits in the format like “0.0.0” where the first digit represents the major version, the second digit represents the minor version, and the last digit represents the patch version. To simulate the evolution process of a project, in Figure 4, we start with evo0 (evolving-aware setting) and increment the minor version number by one each time. If there are no more minor versions under that major version, the major version is incremented until the latest version of the project is reached. We do not update the patch version because patch versions generally entail minimal changes to the project.
As shown in Figure 4, the performance of LLMs deviates from reality as the project evolves. However, the artificial improvement in model performance induced by the evolving-ignored setting does not exhibit a linear increase but rather a curve indicating an upward trend. Updates in the project sometimes introduce irrelevant project contexts, which can act as noise and impact the performance of LLMs. Additionally, project updates may result in the removal or modification of originally useful contexts, leading to the direct loss of context that aids in model objective generation, thereby causing a decline in model performance. These factors may be the primary reasons for the occasional slight decrease in model performance when using different versions of the project as context sources.
RQ2 Summary: In the evolving-ignored setting, as the project evolves, the model’s performance gradually deviates from reality, and the model’s overestimation does not increase linearly but rather follows a fluctuating curve.
5.3. RQ3: How different code descriptions in the dataset affect code generation performance?
During the benchmark collection process, we notice that some projects tend to write detailed docstrings, while others prefer to describe the function’s functionality using only brief statements. To make our benchmark more diverse and reflective of real-world development scenarios, we manually write both brief and detailed docstrings for each function in HumanEvo. In this experiment, we evaluate the performance of the selected LLMs on two types of manually written docstrings. As shown in Table 6, experimental results indicate that the majority of LLMs perform better on detailed docstrings compared to brief docstrings. This is because detailed docstrings provide a more comprehensive description of the functionality of the target function, some even including the implementation logic of the target function. In addition, detailed docstrings also provide explanations for the input and output of the target function, enabling LLMs to better understand the tasks to be performed. Conversely, brief docstrings only introduce the intention of the target function, which may lead to LLMs struggling to understand user intent in certain complex scenarios, ultimately resulting in decreased performance.
However, the decline in LLMs’ performance is not significant, mainly because we adopted a retrieval-augmented generation approach, retrieving a large amount of context from the project repository through similarity calculation. These contexts contribute to a lengthy prompt. However, when LLMs deal with overly long prompts, their attention gets dispersed, leading to the neglect of some useful information by the model.
RQ3 Summary: Compared to brief docstrings, detailed docstrings provide more beneficial information to LLMs, resulting in better performance on detailed docstrings. However, due to the extensive project context required for completing the repository-level code generation task, the model’s attention gets dispersed. Consequently, the performance improvement on detailed docstrings is not substantial.
Table 6. Model performance under different docstrings styles.DepSCode is short for DeepSeekCoder.
Model Docstring Python Java
EI HumanEvo EI HumanEvo
CodeLlama-7B Detailed 29.0% 22.0% 8.5% 6.0%
Brief 28.5% 22.0% 7.0% 4.5%
CodeLlama-13B Detailed 30.5% 23.0% 9.5% 7.0%
Brief 30.0% 22.5% 8.5% 5.0%
CodeLlama-34B Detailed 31.5% 24.0% 14.5% 10.0%
Brief 31.0% 24.5% 11.5% 7.0%
DepSCode-6.7B Detailed 30.0% 23.0% 13.0% 9.0%
Brief 30.0% 23.0% 10.5% 6.5%
DepSCode-33B Detailed 31.0% 25.0% 15.5% 11.5%
Brief 31.0% 24.0% 13.5% 9.0%
GPT-3.5-Turbo Detailed 32.5% 25.5% 18.0% 13.0%
Brief 31.5% 24.5% 14.5% 11.0%
GPT-4 Detailed 34.5% 26.5% 20.5% 14.5%
Brief 32.5% 25.0% 16.5% 12.5%
6. Threads To Validity
First, due to limited computational resources, our experiments did not cover all code LLMs, thus not fully reflecting the performance of all LLMs in actual development scenarios, which may slightly affect the representativeness of our experiments. Second, currently HumanEvo only includes two languages. In the future, we will continue to expand HumanEvo to cover as many programming languages as possible. Third, in prompt format selection, we ultimately chose the one that yielded the best results after trying several formats. However, since we did not cover all available prompt formats, the one we selected may not be the best. Finally, our benchmark includes GitHub data from before 2023, which might already be present in the pretraining dataset of the models, posing a risk of data leakage. To mitigate this risk, we construct an automated benchmark construction pipeline that continuously updates the dataset.
7. Related Work
CoderEval (Yu et al., 2024) contains 230 samples from 43 Python projects and 230 samples from 10 Java projects. They categorize task instances into six runnable-levels to better analyze LLMs’ code generation performance. EvoCodeBench (Li et al., 2024a) collects data from 25 real-world repositories. It also ensures that the distribution of programming tasks (with or without dependencies) matches the distribution of real repositories. RepoCoder (Zhang et al., 2023b) proposed RepoEval, which consists of the latest and high-quality real-world repositories covering scenarios such as line completion, API invocation, and function body completion. Currently, most benchmarks for LLM code generation provide simple functions, mostly targeting general functions rather than specific domains, thus limiting practicality. However, BioCoder (Tang et al., 2024), a benchmark for bioinformatics code generation, has been introduced. BioCoder carefully selects 28 high-quality projects related to bioinformatics, then parses these repositories to generate ASTs and obtain relevant data. A custom code context is written for each filtered function, including necessary imports, cross-file dependencies, and some test cases.
However, all the benchmarks mentioned above have a common flaw: they all overlook the evolution of projects over time. These benchmarks take the latest version of the project repository as the context source, and such an evolving-ignored setting makes it difficult for them to accurately reflect the performance of LLMs in real-world development scenarios. Therefore, in this work, we propose HumanEvo, the first repository-level evolving-aware code generation benchmark. Before obtaining the project context for the target function, we roll back the project repository to the state before the target function was committed to ensure that the project context provided to LLMs does not suffer from issues such as future context leakage and missing useful context. Our objective is to offer LLMs a more realistic development scenario, thus more accurately reflecting the actual performance of LLMs in real-world development scenarios.
8. Conclusion
In this paper, we highlight a previously overlooked flaw in evaluating existing repository-level code generation methods: the evolution of projects over time. To address it, we introduce HumanEvo, an evolving-aware repository code generation benchmark, and conduct a comprehensive empirical study to explore LLMs’ performance in pragmatic development scenarios. Experimental results demonstrate that the evolving-ignored situation leads to the inflated performance of LLMs. By incorporating temporal considerations into the benchmark, HumanEvo provides a more comprehensive assessment of model performance in pragmatic development scenarios. We believe that our empirical results and HumanEvo will offer valuable insights for the future development of LLMs-based repository-level code generation benchmarks.